# E-SQL: Direct Schema Linking via Question Enrichment in Text-to-SQL

### v1

@@ Abstract.
Translating Natural Language Queries into Structured Query Language (Text-to-SQL or NLQ-to-SQL) is a critical task extensively studied by both the natural language processing and database communities, aimed at providing a natural language interface to databases (NLIDB) and lowering the barrier for non-experts. Despite recent advancements made through the use of Large Language Models (LLMs), significant challenges remain. These include handling complex database schemas, resolving ambiguity in user queries, and generating SQL queries with intricate structures that accurately reflect the user’s intent. In this work, we introduce E-SQL, a novel pipeline specifically designed to address these challenges through direct schema linking and candidate predicate augmentation. E-SQL enhances the natural language query by incorporating relevant database items (i.e., tables, columns, and values) and conditions directly into the question, bridging the gap between the query and the database structure. The pipeline leverages candidate predicate augmentation to mitigate erroneous or incomplete predicates in generated SQLs. We further investigate the impact of schema filtering, a technique widely explored in previous work, and demonstrate its diminishing returns when applied alongside advanced large language models. Comprehensive evaluations on the BIRD benchmark illustrate that E-SQL achieves competitive performance, particularly excelling in complex queries with a 66.29% execution accuracy on the test set. All code required to reproduce the reported results is publicly available on our GitHub repository1.
11footnotetext: https://github.com/HasanAlpCaferoglu/E-SQL
@@ 1. Introduction
The task of translating natural language queries into SQL (Text-to-SQL) has garnered considerable attention due to its potential to lower the technical barrier for non-experts and enhance the efficiency of querying or recommendation systems. Situated at the intersection of natural language processing (NLP) and database management, this task aims to enable users to interact with databases through simple, natural language queries, without requiring extensive knowledge of SQL syntax or database schema structures. Despite advancements in leveraging large language models (LLMs) for Text-to-SQL, a significant performance gap of approximately 20% still remains between the best-performing models and human-level accuracy, underscoring that even the most sophisticated pipelines are not yet suitable for real-world deployment as a natural language interface to databases (Li et al., 2024a).
@@
Prior to the emerge of large language models (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a, b; AI, 2024; Chen et al., 2021; Chowdhery et al., 2023; Anil et al., 2023; Li et al., 2023a), works (Zhong et al., 2017; Guo et al., 2019; Lin et al., 2020; Fu et al., 2023; Choi et al., 2021; Wang et al., 2020; Usta et al., 2021, 2023; Lei et al., 2020) focused on building encoder-decoder based neural network architectures leveraging recurrent neural networks  (Hochreiter and Schmidhuber, 1997; Cho et al., 2014) and various pre-trained language models (Devlin et al., 2019; Raffel et al., 2020; Yu et al., 2021; Clark et al., 2020). These early approaches established a foundation but were often limited in handling complex queries or schemas.
@@
Large Language Models (LLMs) have shown substantial potential in the Text-to-SQL task, yielding impressive results across various benchmarks. To further enhance the reasoning abilities of LLMs, a variety of in-context learning (ICL) techniques have been introduced, including chain-of-thought (CoT) prompting (Wei et al., 2024), question decomposition (Zhou et al., 2023; Khot et al., 2023), self-consistency (Wang et al., 2023), and others  (Xi et al., 2023; Huang et al., 2023; Madaan et al., 2023; Zhang et al., 2023). Although many of these strategies have been successfully applied to Text-to-SQL translation pipelines (Pourreza and Rafiei, 2023; Gao et al., 2024; Qu et al., 2024; Li et al., 2024b; Talaei et al., 2024), improving LLM reasoning specifically from the perspective of question refinement remains relatively underexplored.
@@
Beyond in-context learning, LLM performance can also be improved through fine-tuning or training from scratch. However, these techniques are resource-intensive, requiring significant computational resources and large volumes of task-specific annotated data. While proprietary models are less frequently fine-tuned for Text-to-SQL (Maamari et al., 2024), promising results have been achieved through the fine-tuning of numerous open-source models (Pourreza and Rafiei, 2024; Gao et al., 2024; Pourreza and Rafiei, 2023; Li et al., 2024b; Talaei et al., 2024; Wang et al., 2024).
@@
Figure 1 demonstrates the general pipeline and modules used in prior works. Another critical aspect of the Text-to-SQL task is schema linking, as seen in the Figure 1, which involves connecting the sense of natural language query to the database schema. Despite various approaches to enhance schema linking, it remains a core challenge. Schema pruning, a technique commonly used to eliminate irrelevant database items, has been widely adopted to reduce noise for downstream tasks. While both neural network-based schema filtering techniques (Li et al., 2023c, 2024b) and LLM-based schema filtering methods  (Dong et al., 2023; Talaei et al., 2024; Lee et al., 2024; Pourreza and Rafiei, 2023; Qu et al., 2024) have been explored, our findings align with those of Maamari et al. (Maamari et al., 2024), indicating that schema filtering can lead to performance degradation when latest generation LLMs are employed. Furthermore, several studies (Gao et al., 2024; Pourreza and Rafiei, 2023; Talaei et al., 2024; Qu et al., 2024; Wang et al., 2024; Pourreza and Rafiei, 2024; Maamari et al., 2024) reveal that providing database-related information in response to a query significantly enhances performance.
@@
In this work, we introduce E-SQL: Direct Schema Linking via Question Enrichment in Text-to-SQL, a novel pipeline designed to directly address the schema linking challenge through question enrichment and candidate predicate augmentation. We explore the perspective of both LLM reasoning and schema linking improvement through question enrichment. E-SQL enhances the natural language query representation by incorporating relevant database elements—such as tables, columns, and values—directly into the query. This approach is augmented by generating candidate predicates, which reduce the likelihood of erroneous or incomplete SQL predicates. This methodology differs from traditional schema filtering techniques, which have been commonly used to simplify the schema presented to the model. The E-SQL pipeline consists of four main modules: Candidate SQL Generation (CSG), Candidate Predicate Generation (CPG), Question Enrichment (QE), and SQL Refinement (SR).
@@
In the Candidate SQL Generation (CSG) module, an initial SQL query is generated. This query is then parsed to extract values and operations from its predicates. The Candidate Predicate Generation (CPG) module uses these extracted elements to find similar values from the database and constructs candidate predicates. Using the candidate predicates, the Question Enrichment (QE) module instructs the LLM to incorporate relevant database items and possible predicates into the natural language question. Simultaneously, the candidate SQL query is executed to identify potential execution errors. Finally, in the SQL Refinement (SR) module, the candidate SQL query is either refined or a new SQL query is generated, utilizing the enriched question, candidate predicates, and any identified execution errors.
@@
We also explore the efficacy of schema filtering, a widely adopted technique in previous research. We incorporate an additional schema filtering module into our pipeline, where the LLM is instructed to select only the database tables and columns relevant to the query while eliminating others. Following this, the Filtered Schema Correction technique is applied to resolve any inconsistencies between the filtered schema and the original database schema. Our experiments demonstrate that schema filtering can negatively affect performance when applied in conjunction with advanced LLMs. Instead, direct schema linking through question enrichment and candidate predicate augmentation provides a more reliable strategy for accurate SQL generation, particularly in complex cases.
@@
Through an ablation study detailed in Section 4.4, we illustrate the effectiveness of each module within the pipeline for the Text-to-SQL task. In particular, our question enrichment module significantly improves performance on challenging questions, yielding nearly a 5% increase in accuracy.
@@
We evaluate E-SQL on the BIRD benchmark (Li et al., 2024a), a well-known standard for the Text-to-SQL task, and demonstrate its ability to handle complex queries while maintaining competitive performance with state-of-the-art methods. Our findings suggest that the integration of enriched questions and candidate predicates leads to more accurate SQL generation, particularly for complex queries involving multiple conditions and joins. Therefore, our approach establishes a new paradigm for schema linking and prompt augmentation in the context of Text-to-SQL translation.
Our key contributions to the Text-to-SQL task, which address both schema linking and LLM reasoning capabilities, can be summarized as follows:
@@
  • We propose a new paradigm for schema linking through question enrichment, which leads to direct schema linking by incorporating related database items and potential conditions into the natural language question.
@@
  • To the best of our knowledge, we are the first to enhance both LLM reasoning capabilities and schema linking performance through the use of a question enrichment module in Text-to-SQL translation task.
@@
  • Our experiments also confirm the potential drawbacks of traditional schema filtering techniques, particularly when combined with advanced large language models.
@@ 2. Related Work
2.1. Supervised Fine-Tuned (SFT) Neural Networks in Text-to-SQL
Prior to the introduction of GPT-3 (Brown et al., 2020), a large-scale pre-trained autoregressive language model, the primary approaches to the Text-to-SQL translation task were based on encoder-decoder neural network architectures.
In the Text-to-SQL translation task, encoding the question and database components was a critical step for high translation performance. Encoding the users’ natural language question serialized with database items allows neural networks to create a relation between question and database schema, called schema linking. The serialized input is encoded with bidirectional LSTM (Hochreiter and Schmidhuber, 1997; Schuster and Paliwal, 1997) in Seq2SQL (Zhong et al., 2017) and IRNet (Guo et al., 2019), while pre-trained language models (PLM) such as BERT (Devlin et al., 2019), T5 (Raffel et al., 2020), GraPPa (Yu et al., 2021) and Electra (Clark et al., 2020) were utilized later to produce high quality contextualised representations. BRIDGE (Lin et al., 2020) and CatSQL (Fu et al., 2023) used BERT and GraPPa, respectively. Differently, convolutional neural networks were utilized in RYANSQL (Choi et al., 2021), while graph neural networks were utilized in RATSQL (Wang et al., 2020) to encode the input. For finer schema linking, Usta et al. (Usta et al., 2021, 2023) propose DBTagger for the keyword mapping problem, constructing a mapping between tokens in the query and relational database elements, by utilizing multi-task learning leveraging part-of-speech (POS) tags of keywords. SLSQL proposed by Lei et al. (Lei et al., 2020) provides a complete end-to-end solution for the Text-to-SQL task leveraging keyword mapping. Inspired by DBTagger and SLSQL, in the question enrichment phase, we utilize schema tags injecting into the question.
On the decoding side, the process of generating SQL queries from the encoded input has been handled in several ways. The most straightforward method considers the generation of SQL as a sequence generation task, a technique employed by several Text-to-SQL models such as Seq2SQL and BRIDGE (Zhong et al., 2017; Lin et al., 2020). An alternative grammar-based approach generates an abstract syntax tree (AST) (Yin and Neubig, 2017), which is considered as intermediate representation of SQL and shows higher level abstraction and used in models such as IRNet (Guo et al., 2019), RYANSQL (Choi et al., 2021), SLSQL (Lei et al., 2020) and RATSQL (Wang et al., 2020). Another effective strategy is the sketch-based slot-filling approach, where the neural network does not need to generate SQL keywords but rather fills predefined slots in a sketch-like structure (Fu et al., 2023; Xu et al., 2017; Yu et al., 2018).
2.2. LLMs in Text-to-SQL
With the emergence of both proprietary and open-source Large Language Models (LLMs) such as GPT-3, instructGPT, GPT-4, LLama family, Codex, PaLM, StarCoder CodeLlama and DeepSeek-Coder (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a, b; AI, 2024; Chen et al., 2021; Chowdhery et al., 2023; Anil et al., 2023; Li et al., 2023a; Guo et al., 2024; Rozière et al., 2024) has marked a significant shift in the field. Thanks to their advanced reasoning and comprehension capabilities, the research community has increasingly focused on harnessing the power of these models for Text-to-SQL tasks.
2.2.1. LLM Reasoning
@@
A breakthrough method proposed by Wei et al. (Wei et al., 2024) improves the ability of large language models on complex and multi-step reasoning tasks significantly by leading them to generate a series of intermediate reasoning steps, chain-of-thought(CoT). Kojima et al. (Kojima et al., 2022) explored a simple yet effective prompting, ”Let’s think step by step”, to uncover the reasoning ability of the language models in zero-shot mode. Zhou et al. (Zhou et al., 2023), Khot et al. (Khot et al., 2023) and Dua et al. (Dua et al., 2022) leverage the idea of decomposing complex questions into several simpler sub-questions and constructing a solution by generating an answer for each sub-question. Self-Consistency, proposed by Wang et al. (Wang et al., 2023), utilizes the majority voting technique to select the most consistent answer over a set of answers generated by diverse reasoning paths. Huang et al. proposed Self-Improve (Huang et al., 2023), a technique that enhances the reasoning capabilities of large language models by self-training on self-generated data, utilizing Chain-of-Thought (CoT) prompting, self-consistency, and supervised labels. Self-Refine (Madaan et al., 2023) demonstrates how to generate accurate and consistent answers for complex tasks by iteratively refining the language model’s responses through a self-feedback mechanism, using only a single LLM. While previous efforts primarily focused on the answer generation stage to enhance the reasoning capabilities of language models, the Self-Polish technique proposed by Xi et al. (Xi et al., 2023) demonstrates significant improvements by refining problem contexts and questions. Inspired by Self-Polish, our work aims to generate clear, well-focused, and schema-aware natural language queries, free from irrelevant information.
2.2.2. Schema Linking and Filtering
@@
Previous works have shown that removing irrelevant database elements can improve schema linking and enhance model performance in Text-to-SQL tasks by reducing the likelihood of schema-based hallucinations. This process, known as schema filtering or schema pruning, has been extensively studied. RESDSQL (Li et al., 2023c) addresses this by classifying schema items based on their relevance to the natural language query and then filtering them according to their classification probabilities after ranking. Li et al. (Li et al., 2024b) follow the strategy applied in RESDSQL for schema filtering. While C3 (Dong et al., 2023), CHESS (Talaei et al., 2024), and MCS-SQL (Lee et al., 2024) first filter database tables (table linking) and then select the most appropriate columns (column linking) from the previously filtered tables, DIN-SQL (Pourreza and Rafiei, 2023) uses a single step to select only the question relevant database tables and columns. For schema linking and filtering, the TASL module of TA-SQL, proposed by Qu et al. (Qu et al., 2024), generates a list of schema entities (tables and columns) from an LLM-generated dummy SQL query, which is then used to create symbolic representations. Our approach also leverages the generated SQL query; however, rather than using it for schema linking and filtering, we extract possible database values from conditions to generate potential predicates, which are then used in downstream tasks as part of data augmentation. Gao et al. (Gao et al., 2024) explored the impact of different schema representations and demonstrated that representing the database schema as code, instead of natural language representation, leads to a better understanding of the question-to-SQL mapping, thereby improving schema linking. While previous works have demonstrated the positive impact of schema filtering on schema linking, Maamari et al. (Maamari et al., 2024) argue that explicit schema filtering remains useful for less capable large language models (LLMs), but becomes unnecessary with the latest, more advanced LLMs such as GPT-4o. Our experiments on the Text-to-SQL task corroborate this finding, showing that explicit schema filtering is redundant with more sophisticated LLMs.
2.2.3. Data Augmentation

### v2

@@ Abstract.
Translating Natural Language Queries into Structured Query Language (Text-to-SQL or NLQ-to-SQL) is a critical task extensively studied by both the natural language processing and database communities, aimed at providing a natural language interface to databases (NLIDB) and lowering the barrier for non-experts. Despite recent advancements made through the use of Large Language Models (LLMs), significant challenges remain. These include handling complex database schemas, resolving ambiguity in user queries, and generating SQL queries with intricate structures that accurately reflect the user’s intent. In this work, we introduce E-SQL, a novel pipeline specifically designed to address these challenges through direct schema linking and candidate predicate augmentation. E-SQL enhances the natural language query by incorporating relevant database items (i.e., tables, columns, and values) and conditions directly into the question and SQL construction plan, bridging the gap between the query and the database structure. The pipeline leverages candidate predicate augmentation to mitigate erroneous or incomplete predicates in generated SQLs. Comprehensive evaluations on the BIRD benchmark illustrate that E-SQL achieves competitive performance, particularly excelling in complex queries with a 66.29% execution accuracy on the test set. A further observation from our experiments reveals that incorporating schema filtering into the translation pipeline does not have a positive impact on performance when the most advanced proprietary LLMs are used. Additionally, our experiments with small LLMs highlight the importance and positive impact of enriched questions on their performance. Without fine-tuning, single-prompt SQL generation using enriched questions with DeepSeek Coder 7B Instruct 1.5v achieves 56.45% execution accuracy on the BIRD development set.
Text-to-SQL, Large Language Model, Schema Linking, Question Enrichment
@@ 1. Introduction
The task of translating natural language queries into SQL (Text-to-SQL) has garnered considerable attention due to its potential to lower the technical barrier for non-experts and enhance the performance of querying or recommendation systems. Situated at the intersection of natural language processing (NLP) and database management, this task aims to enable users to interact with databases through simple, natural language queries, without requiring extensive knowledge of SQL syntax or database schema structures. Despite advancements in utilizing large language models (LLMs) for Text-to-SQL, a significant performance gap of around 20% still remains between the best-performing models and human-level accuracy, underscoring that even the most sophisticated pipelines are not yet suitable for real-world deployment as a natural language interface to databases (Li et al., 2024a).
@@
Prior to the emerge of LLMs (Brown et al., 2020; Ouyang et al., 2022; OpenAI, 2023; Touvron et al., 2023a, b; AI, 2024; Chen et al., 2021; Chowdhery et al., 2023; Anil et al., 2023; Li et al., 2023a), a wide range of studies  (Zhong et al., 2017; Guo et al., 2019; Lin et al., 2020; Fu et al., 2023; Choi et al., 2021; Wang et al., 2020; Usta et al., 2021, 2023; Lei et al., 2020) focused on building encoder-decoder based neural network architectures utilizing recurrent neural networks  (Hochreiter and Schmidhuber, 1997; Cho et al., 2014) and various pre-trained language models (Devlin et al., 2019; Raffel et al., 2020; Yu et al., 2021; Clark et al., 2020). These early approaches established a foundation but were often limited in handling complex queries or schemas.
@@
LLMs have shown substantial potential in the Text-to-SQL task, yielding impressive results across various benchmarks. To further enhance the reasoning capabilities of LLMs, a variety of in-context learning (ICL) techniques have been introduced, including chain-of-thought (CoT) prompting (Wei et al., 2024), question decomposition (Zhou et al., 2023; Khot et al., 2023), self-consistency (Wang et al., 2023), and others  (Xi et al., 2023; Huang et al., 2023; Madaan et al., 2023; Zhang et al., 2023). Although many of these strategies have been successfully applied to Text-to-SQL translation pipelines (Pourreza and Rafiei, 2023; Gao et al., 2024; Qu et al., 2024; Li et al., 2024c; Talaei et al., 2024), improving LLM reasoning specifically from the perspective of question refinement remains relatively underexplored.
@@
Beyond in-context learning, LLM performance can also be improved through fine-tuning or training from scratch. However, these techniques are resource-intensive, requiring significant computational resources and large volumes of task-specific annotated data. While proprietary models are less frequently fine-tuned for Text-to-SQL (Maamari et al., 2024), promising results have been achieved through the fine-tuning of numerous open-source models (Pourreza and Rafiei, 2024; Gao et al., 2024; Pourreza and Rafiei, 2023; Li et al., 2024c; Talaei et al., 2024; Wang et al., 2024).
@@
Figure 1 demonstrates the general pipeline and modules used in prior works. A critical component of the Text-to-SQL task is schema linking, which involves connecting the sense of natural language query to the database schema. Although various methods have been proposed to enhance schema linking, it remains a core challenge. Schema filtering, a technique commonly used to eliminate irrelevant database items, has been widely adopted to reduce noise for downstream tasks. While both neural network-based (Li et al., 2023c, 2024c) and LLM-based (Dong et al., 2023; Talaei et al., 2024; Lee et al., 2024; Pourreza and Rafiei, 2023; Qu et al., 2024) schema filtering techniques have been explored, our findings align with those of Maamari et al. (Maamari et al., 2024), indicating that schema filtering can result in performance degradation when the latest generation LLMs are employed. Additionally, several studies (Gao et al., 2024; Pourreza and Rafiei, 2023; Talaei et al., 2024; Qu et al., 2024; Wang et al., 2024; Pourreza and Rafiei, 2024; Maamari et al., 2024) reveal that providing database-related information in response to a query significantly enhances performance.
@@
In this work, we introduce E-SQL: Direct Schema Linking via Question Enrichment in Text-to-SQL [^1] The complete code required to reproduce the reported results is publicly available on our GitHub repositoryhttps://anonymous.4open.science/r/E-SQL_Direct_Schema_Linking, a novel pipeline designed to directly address the schema linking challenge through question enrichment and candidate predicate augmentation. We explore the improvement of both LLM reasoning and schema linking from the perspective of question enrichment. E-SQL enhances the natural language query representation by incorporating relevant database elements—such as tables, columns, and values—directly into the query, along with an associated SQL construction plan. This approach is augmented by generating candidate predicates, which reduce the likelihood of erroneous or incomplete SQL predicates. This methodology differs from traditional schema filtering techniques, which have been commonly used to simplify the schema presented to the model. The E-SQL pipeline consists of four main modules: Candidate SQL Generation (CSG), Candidate Predicate Generation (CPG), Question Enrichment (QE), and SQL Refinement (SR).
@@
In the Candidate SQL Generation (CSG) module, an initial SQL query is generated. This query is then parsed to extract values and operations from its predicates. The Candidate Predicate Generation (CPG) module uses these extracted elements to find similar values from the database and constructs candidate predicates. Using the candidate predicates, the Question Enrichment (QE) module instructs the LLM to incorporate relevant database items and possible predicates into the natural language question, while concurrently formulating SQL construction steps as its reasoning process. These steps are then utilized to produce a fully enriched query. Simultaneously, the candidate SQL query is executed to identify potential execution errors. Finally, in the SQL Refinement (SR) module, the candidate SQL query is either refined or a new SQL query is generated, utilizing the enriched question, candidate predicates, and any identified execution errors.
@@
The impact of schema filtering, a widely adopted technique in previous research, is also explored on our pipeline. We incorporate an additional schema filtering module into our pipeline, where the LLM is instructed to select only the database tables and columns relevant to the query while eliminating others. Following this, the Filtered Schema Correction technique is applied to resolve any inconsistencies between the filtered schema and the original database schema. Our experiments demonstrate that schema filtering can negatively affect performance when applied in conjunction with the most advanced proprietary LLMs. Instead, direct schema linking through question enrichment and candidate predicate augmentation provides a more reliable strategy for accurate SQL generation, particularly in complex cases.
@@
Through an ablation study, we illustrate the effectiveness of each module within the pipeline for the Text-to-SQL task. In particular, our question enrichment module significantly improves performance on challenging questions, yielding nearly a 5% increase in accuracy.
@@
We evaluate E-SQL on the Spider (Yu et al., 2018) and BIRD (Li et al., 2024a) benchmarks, well-known standard datasets for the Text-to-SQL task, and demonstrate its ability to handle complex queries while maintaining competitive performance with state-of-the-art methods. Our findings suggest that the integration of enriched questions, SQL generation steps, and candidate predicates leads to more accurate SQL generation, particularly for complex queries involving multiple conditions and joins. Therefore, our approach establishes a new paradigm for schema linking and prompt augmentation by leveraging question enrichment and candidate predicate augmentation in the context of Text-to-SQL translation.
The key contributions of our work can be summarized as follows:
@@
  • We propose a new paradigm for schema linking through question enrichment, which leads to direct schema linking by incorporating related database items and potential conditions into the natural language question. Fully enriched queries further guide LLMs in SQL construction by providing explicit logical steps.
@@
  • To the best of our knowledge, we are the first to enhance both LLM reasoning capabilities and schema linking performance through the use of a question enrichment module, which composes database integrated questions, in Text-to-SQL translation task.
@@
  • Our experiments also confirm the potential drawbacks of traditional schema filtering techniques when integrated into a Text-to-SQL translation pipeline like ours, which leverages the most advanced proprietary LLMs.
  • We demonstrate the importance and positive impact of database-integrated questions, including logical SQL construction steps, on the performance of small LLMs in the task of Text-to-SQL translation, achieved without requiring fine-tuning.
@@ 2. Related Work
Before the emergence of LLMs, supervised fine-tuning approaches in Text-to-SQL translation focused on encoder - decoder architectures that utilized recurrent neural networks (RNNs)  (Zhong et al., 2017; Guo et al., 2019; Usta et al., 2021), pre-trained language models (PLMs)  (Lin et al., 2020; Fu et al., 2023; Lei et al., 2020), convolutional neural networks (CNNs)  (Choi et al., 2021) and graph neural networks (GNNs) (Wang et al., 2020). These methods encoded natural language questions alongside database schema to establish schema linking and generated SQL queries through sequence generation  (Zhong et al., 2017; Lin et al., 2020), grammar-based methods  (Guo et al., 2019; Choi et al., 2021; Wang et al., 2020), or sketch-based slot-filling strategies. These approaches provided a foundational understanding for Text-to-SQL tasks, paving the way for more advanced solutions.
The emergence of both proprietary and open-source LLMs (Brown et al., 2020; Ouyang et al., 2022; Touvron et al., 2023a; Chen et al., 2021; Hui et al., 2024; Guo et al., 2024) has marked a significant shift in the field. Thanks to their advanced reasoning and comprehension capabilities, the research community has increasingly focused on harnessing the power of these models for Text-to-SQL tasks.
2.1. LLM Reasoning
@@
The Chain-of-Thought (CoT) approach (Wei et al., 2024) significantly improves the performance for multi-step reasoning tasks by guiding LLMs to generate intermediate reasoning steps. Kojima et al. (Kojima et al., 2022) explored a simple yet effective prompting, ”Let’s think step by step”, to uncover the reasoning ability of the language models in zero-shot mode. Decomposing complex questions into simpler sub-questions has been explored (Zhou et al., 2023; Khot et al., 2023; Dua et al., 2022). Techniques like Self Consistency (Wang et al., 2023) apply majority voting to select consistent answers, while Self-Improve (Huang et al., 2023) and Self-Refine (Madaan et al., 2023) iteratively refine responses through self-generated data and feedback, respectively. Moving beyond answer generation, Xi et al. (Xi et al., 2023) demonstrated significant gains by refining problem contexts and questions through the Self-Polish technique.
To the best of our knowledge, the refinement and generation of high-quality user queries, particularly with embedded reasoning that facilitates the construction of correct SQL queries, has not been explored in the literature for the Text-to-SQL translation task. Our work addresses this gap by proposing a methodology that produces clear, schema-aware queries enriched with reasoning elements, explicitly guiding the generation of accurate SQL queries and overcoming schema linking challenges.
2.2. Schema Linking and Filtering
@@
Previous works have shown that removing irrelevant database elements can improve schema linking and enhance model performance in Text-to-SQL tasks by reducing the likelihood of schema-based hallucinations. This process, known as schema filtering or schema pruning, has been extensively studied. RESDSQL (Li et al., 2023c) and CodeS (Li et al., 2024c) address this by classifying schema items based on their relevance to the natural language query and then filtering them according to their classification probabilities after ranking. While DIN-SQL (Pourreza and Rafiei, 2023) uses a single step to select only the question relevant database tables and columns, C3 (Dong et al., 2023), CHESS (Talaei et al., 2024), and MCS-SQL (Lee et al., 2024) first filter database tables (table linking) and then select the most appropriate columns (column linking) from the previously filtered tables, achieving better schema filtering. For schema linking and filtering, the TASL module of TA-SQL (Qu et al., 2024) generates a list of schema entities (tables and columns) from an LLM-generated dummy SQL query, which is then used to create symbolic representations. Our approach also leverages the initially generated candidate SQL query; however, rather than using it for schema filtering, we extract possible database values from conditions to generate potential predicates, which are then used in downstream tasks as part of data augmentation, explained in Section 3.3. Gao et al. (Gao et al., 2024) explored the impact of different schema representations and demonstrated that representing the database schema as code, instead of natural language representation, leads to a better understanding of the question-to-SQL mapping, thereby improving schema linking. While previous works have demonstrated the positive impact of schema filtering on schema linking, Maamari et al. (Maamari et al., 2024) argue that explicit schema filtering remains useful for less capable large language models (LLMs), but it becomes unnecessary with the latest, more advanced LLMs such as GPT-4o. Our experiments with single-step schema filtering implemented in the E-SQL pipeline corroborate this finding, indicating that explicit schema filtering can be redundant and can result in performance degradation when applied within pipelines leveraging advanced LLMs such as GPT-4o and GPT-4o-mini  (OpenAI, 2023).
2.3. Data Augmentation